OpenStudy (anonymous):

A silverwire has a resistance of 1.5Ω at 0°C and a temperature coefficient of resistance of 0.00375/°C. To what temperature must be the raised to double the resistance?

OpenStudy (anonymous):

I had a smiliar problem expect I was asked to find the R. I used R= (Delta V ^2 x time)/mc(T2-T1)

OpenStudy (anonymous):

so that is the formula?

OpenStudy (anonymous):

My bad, that's not the correct formula for this type of problem. You aren't given the potential difference.

OpenStudy (anonymous):

:( actually i can't understand the problem

OpenStudy (kropot72):

\[R _{t}=R _{0}+\alpha t\] \[R _{t}-R _{0}=\alpha t\] 3 - 1.5 = 0.00375 * t Now solve for t

OpenStudy (anonymous):

why 3?

OpenStudy (kropot72):

Raising the temperature needs to double the resistance. 1.5 ohms * 2 = 3 ohms

OpenStudy (anonymous):

t= 3-15 / 0.0037 ??

OpenStudy (kropot72):

\[t=\frac{3-1.5}{0.00375}\]

OpenStudy (anonymous):

thanks. so the unit must be degree celsius ?

OpenStudy (kropot72):

Yes. The unit for the answer is degrees Celsius :)

OpenStudy (anonymous):

wow thanks a lot. :D

OpenStudy (kropot72):

You're welcome :)